OpenStudy (anonymous):

Would someone please help me see how to work out the problem f/g, when f(x) = square root of 3-x, and g(x) = square root of x^2-1 please? and I need to figure out the domain. Thank you

OpenStudy (klimenkov):

\(\frac{f(x)}{g(x)}=\frac{\sqrt{3-x}}{\sqrt{x^2-1}}\). Domain: \(\left\{\begin{matrix} 3-x\geqslant0\\ x^2-1>0 \end{matrix}\right.\), \(\left\{\begin{matrix} x\leqslant3\\ x\in(-\infty,-1)\bigcup(1,+\infty) \end{matrix}\right.\) \(\Rightarrow x\in(-\infty,-1)\bigcup(1,3]\)